# Local Differential Privacy for Tensors in Distributed Computing Systems

### v2

@@ Abstract
Tensor-valued data, increasingly common in distributed big data applications like autonomous driving and smart healthcare, poses unique challenges for privacy protection due to its multidimensional structure and the risk of losing critical structural information. Traditional local differential privacy methods, designed for scalars and matrices, are insufficient for tensors, as they fail to preserve essential relationships among tensor elements. We introduce TLDP, a novel LDP algorithm for Tensors, which employs a randomized response mechanism to perturb tensor components while maintaining structural integrity. To strike a better balance between utility and privacy, we incorporate a weight matrix that selectively protects sensitive regions. Both theoretical analysis and empirical findings from real-world datasets show that TLDP achieves superior utility while preserving privacy, making it a robust solution for high-dimensional tensor data.
Index Terms:
Local differential privacy, tensor privacy, randomized response
I Introduction
High dimensional tensor data are passionately required in a diverse array of distributed big data applications, such as autonomous driving [1] and smart healthcare [2]. In these systems, tensor data that contains sensitive private information, such as location data, facial images, or medical records, is frequently transmitted between different parties, which creates a significant risk of privacy leakage. For example, through the intercepted tensor-valued features or model parameters (in federated systems), an attacker can easily infer whether a data type was used for training the model. Typically, Local Differential Privacy (LDP) constrains an attacker’s ability to infer details about individual data from the shared data.
@@ I Introduction
Developing schemes for tensors to achieve local differential privacy is very challenging due to the complexity of high-dimensional data. Generally, the noise introduced to the raw data is deeply affected by the sensitivity of the query function, one key factor of which is the size of the tensor. Essentially, the larger the specifications of the data tensors, the more substantial the noise tends to be, potentially leading to a decrease in the accuracy of the results. Extensive research explored LDP schemes for scalars [3, 4, 5], vectors [6, 7], and matrices [8, 9] by either adding random noise or randomly perturbing individual inputs to achieve local differential privacy and achieved promising results, but an LDP approach applicable to tensor-valued data in distributed computing systems has yet to be established. Although a few research studies like [10] introduce an LDP for tensors for distribution estimation, they cannot be directly applied for locally differential privacy protection in distributed computing systems. It is because although the perturbed tensor-valued data still retain certain statistical properties, the model analysis results on such data will be seriously affected. Therefore, in this work, we aim to tackle this challenge by developing a local differential privacy mechanism that ensures the effectiveness of the subsequent data analytics despite the high-dimensionality of tensors.
@@
Both privacy and utility play a critical role when practitioners apply LDP in practical distributed computing systems. However, designing an LDP that balances the privacy and utility of data presents a significant challenge due to the inherent trade-off between the two. This issue is even more pronounced with tensors, where excessive noise/perturbation is usually added due to the high dimensionality of tensors, which can severely diminish their usefulness. To mitigate such degradation, various methods have been explored, as seen in [11, 12, 13]. However, most of them are for scalars or vectors instead of tensors, lacking a utility guarantee for tensors in distributed computing systems. In contrast, our approach ensures practical utility while satisfying $\epsilon-$LDP in distributed computing systems.
@@
In this paper, we formalize the definition of LDP for tensor-valued data and innovatively introduce a Local Differential Privacy algorithm designed for Tensors of varying dimensions named TLDP, preserving the tensor-valued data’s structural information. TLDP incorporates locally differentially private and tensor-shaped noise into the tensors by randomly selecting components of the tensor with a randomized response mechanism to accomplish local differential privacy. The key idea behind it is to randomly choose whether or not to add Laplace/Gaussian noise to the tensor data so that the total amount of noise added to the tensors is lower than directly adding noise to each entity of the tensor without breaking the condition of satisfying $\epsilon-$LDP. We rigorously prove that TLDP satisfies local differential privacy both when using the Laplace/Gaussian noise and has a lower expected error than previous studies.
@@
Additionally, we observe that different parts of a tensor may have varying sensitivities in practical applications. For example, in a road scene image, the facial or license plate number information is more critical than the background for scene understanding tasks in autonomous driving applications. The data privacy of such tasks can be further improved by incorporating a precisely tailored weight matrix that indicates the areas of the tensor that are more important than others. Therefore, we design an optimized local differential privacy mechanism based on TLDP by applying varying weights to different parts of the tensor.
@@
  • We propose an algorithm that ensures $\epsilon-$LDP for tensor-valued data named TLDP. TLDP has a lower expected error and greater utility compared to prior methods, independent of the tensor’s shape. To the best of our knowledge, it is the first work that addresses local differential privacy of tensor-valued data in machine learning-based distributed computing systems.
@@ II Related Work
II-A Primitive algorithms
II-A1 Local differential privacy by adding random noise
Primitive algorithms refer to those that protect data privacy by randomly adding random noise to individual data. They ensure privacy independently and do not rely on any external mechanisms. They include the Gaussian mechanism [14], Laplace mechanism [15], Matrix Variate Gaussian (MVG) [16], and Independent Directional Noise (IDN) [17].
Although the mechanisms mentioned above do not directly apply to tensor-valued queries, our research is nevertheless grounded in the fundamental principle of preserving privacy through the injection of noise. The Gaussian mechanism introduces independent and identically distributed (i.i.d.) Gaussian noise, calibrated according to the $l_{2}$-sensitivity of the query, guarantees $(\epsilon,\delta)$-differential privacy. Similarly, by incorporating noise derived from the Laplace distribution, scaled according to the query function’s $l_{1}$-sensitivity, the Laplace mechanism enforces strong $\epsilon$-differential privacy. The MVG mechanism is tailored for matrix-valued queries and incorporates matrix-valued noise to ensure $(\epsilon,\delta)$-differential privacy. The $l_{2}$-sensitivity of the MVG mechanism is measured using the Frobenius norm, which quantifies the discrepancy between two neighboring matrices. The IDN mechanism achieves differential privacy by appending a noise tensor of equal magnitude to the data tensor. However, LDP demands a higher level of privacy than traditional differential privacy, necessitating more noise to achieve comparable privacy guarantees, which can significantly degrade data utility due to extensive perturbation. Our mechanism is specifically tailored to address the complexities of tensor data. We demonstrate that our mechanism attains $\epsilon$-local differential privacy for tensors, as detailed in Section III. This innovation bridges the gap between traditional local differential privacy mechanisms and the sophisticated requirements of high-dimensional data, preserving privacy while maintaining utility.
Beyond these fundamental mechanisms, several sophisticated approaches have also been introduced. Mechanisms that derive their privacy guarantees from the above mechanisms are well-established, as exemplified by the composition schemes referenced in [18, 19, 20, 21, 9, 7, 6]. Among these, [18] presents an innovative accounting approach for the Gaussian mechanism that minimizes the total additive noise while preserving the same level of privacy protection. In contrast, Kairouz et al. [22] propose an optimal composition scheme that is applicable to a broader range of noise distributions. The studies in [19, 21] provide dynamic accounting methods that adjust based on the algorithm’s runtime convergence. Moreover, [20] ensures a global differential privacy guarantee in environments where datasets are continuously expanding. [9] applies differential privacy to high-order, high-dimensional sparse tensors within the IoT transmission context through individual randomized responses. However, when the central server is required to receive tensor information directly from users, this approach risks disclosing the user’s original information with a certain probability, thereby undermining the purpose of privacy protection. [7] introduced a sparse weight matrix that computes the non-zero elements of the matrix, thereby reducing the computational load of LDP when applied to large-scale datasets. Furthermore, [6] proposed a new privacy budget recovery mechanism, called ChainDP, which adds noise sequentially, so that the noise added by the previous user can be used by subsequent users, thereby reducing the overall noise level and estimation error.
II-A2 Local differential privacy by randomized response to a query
@@
Apart from adding noise to individual data, random response is another popular approach to achieve local differential privacy. Stanley L. Warner et al. [3] introduced the randomized response mechanism in 1965, which is a technique specifically designed to preserve respondent privacy, commonly referred to as $w$-RR. The key idea is to introduce randomness into the response process, allowing respondents to answer sensitive questions without fear of exposure, while still enabling researchers to estimate the true distribution of responses. The $w$-RR technique is specifically designed for discrete datasets with exactly two distinct values. To extend its applicability, two primary paths for improvement can be explored. The first involves encoding and transforming the variable’s multiple values, as demonstrated in methods such as Randomized Aggregatable Privacy-Preserving Ordinal Response (RAPPOR) [4] and S-Hist [23], to ensure compatibility with the binary nature expected by $w$-RR. The RAPPOR employs Bloom filters and hash functions to encode a single element into a vector, then perturb the original value through both permanent and instantaneous randomized responses. It represents a method for collecting statistical data from end-user applications while preserving anonymity and ensuring robust privacy protections. However, RAPPOR has two main drawbacks: (1) the transmission overhead between the user and the data aggregator is relatively high, as each user is required to send a vector whose length is determined by the size of its bloom filter; (2) the data collector must pre-collect a list of candidate strings for frequency counting.
@@
To address the first issue, the S-Hist method [23] offers a solution in which each user encodes the string, randomly selects a bit, applies random response techniques to perturb it, and then sends it to the data collector. This approach significantly reduces the transmission cost by minimizing the amount of data each user needs to send. For the second problem, building upon the RAPPOR-based encoding-decoding framework, Kairouz et al. [24] further proposed the $O$-RR method by introducing hash mapping and grouping operations. Hash mapping allows the method to focus on the encoded values rather than the original strings, eliminating the need to pre-collect a list of candidate strings. Additionally, the use of grouping operations reduces the probability of hash mapping value collisions. This combination of techniques enhances the robustness and flexibility of differential privacy in various data collection scenarios.
To overcome the second limitation of $w$-RR, one approach is to improve the distribution of the $w$-RR technique, making it applicable to variables with more than two values. This is achieved through mechanisms such as $k$-RR [22]. In the $k$-RR method, the original data is preserved with probability $p$ and is flipped to any other value with a probability of $\frac{1-p}{n-1}$, where $n$ is the number of possible values. This mechanism ensures that the perturbation is uniformly distributed across all possible values, thereby preserving the privacy of the data.
The techniques discussed above are designed for discrete values. For continuous values, however, they should first be discretized to meet the fundamental requirement of the $w$-RR technique before privacy protection can be applied. This concept is primarily applied in the context of local differential privacy for mean statistics methods like MeanEst [25, 26] and Harmony-mean [27].
II-B Learning With Differential Privacy
@@
In summary, we consider our paper introduces a primitive local differential privacy scheme, which fits well in distributed computing systems but certainly can be applied in any other machine learning systems, where individual users’ tensor-valued data is private and requires to be preserved.
III Preliminaries
In this section, we aim to provide our readers with foundational knowledge to facilitate the comprehension of our work.
III-A Local Differential Privacy
Local differential privacy (LDP) enables data collection and analysis while ensuring strong privacy protection by perturbing data before it is shared with a central aggregator. This ensures that even the data collector cannot determine an individual’s true value with certainty. The privacy guarantee of LDP is expressed by bounding the logarithmic ratio of the output probabilities for any two possible inputs from a single user, ensuring that the private mechanism $M$ does not reveal too much about the user’s true input. This local and individualized privacy guarantee makes LDP particularly suitable for scenarios where users do not trust the data collector and need to protect their data before sharing it. Formally, given any two tensor-valued inputs from the same user $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, $y$ be the output and $M$ be the private scheme, the following condition holds:
@@ III-A Local Differential Privacy
$\forall y\in Range(\textbf{M}),Pr[\textbf{M}(\mathcal{X}_{1})=y]\leq e^{\epsilon}Pr[\textbf{M}(\mathcal{X}_{2})=y]$
@@
Problem Definition and Notations. We suppose that in this article, each user has a number of tensor-valued data, and $\mathcal{X}\in\mathbf{R}^{I_{1}\times I_{2}\times\dots\times I_{N}}$ is any one of them. We utilize $\Delta$ to represent the range of the values of a tensor in a dataset and $[\Delta]$ to denote all possible values of the tensor. For example, $\Delta=256$ and $[\Delta]=\{0,1,2,\dots,255\}$ for an image of the MNIST dataset because that range of the values of an image of the MNIST dataset is 0-255. Without compromising generality, we assume the input domain is $[\Delta]$, i.e., $\mathcal{X}_{i_{1}i_{2}\dots i_{N}}\in[\Delta]$. If the tensor takes continuous values, it is recommended to discretize it initially. The notations used in this paper are summarized in Table I for easy reference.
TABLE I: Notations
@@
$I$ Product of the tensor’s dimensions, i.e., $I=I_{1}I_{2}\cdots I_{N}$
@@
Our goal is to perturb the original private tensor before sharing it with other parties, ensuring that users’ sensitive information can be properly preserved without compromising data utility. The parameter $\epsilon$, defined as the upper bound of the KL-divergence between the data distributions before and after perturbation, quantifies the effectiveness of privacy-preserving algorithms. As stated in Definition 1, a lower value of $\epsilon$ corresponds to stronger privacy guarantees. To access utility, we evaluate the F1-score of the model parameters trained from the perturbed data as a measure of the data’s effectiveness. Specifically, we train the model with the perturbed data and then validate it with the unperturbed data. The closer the validation results are to the original, the higher the utility. It is evident that as privacy protection increases (i.e., $\epsilon$ decreases), the accuracy of the trained model tends to decline. Our research seeks to strike a balance between data utility and privacy protection.
III-B Randomized Response
Randomized response is a statistical technique introduced in the 1960s by [3] for gathering data on sensitive topics while allowing respondents to maintain confidentiality. Surveys employing this method enable the accurate calculation of precise population statistics while ensuring individual privacy. Today, it is the dominant perturbation mechanism used in local differential privacy protection of scalars. For binary-valued problems, the randomized response method retains the original data with probability $p$ and flips it with probability $1-p$. In most cases, $p>\frac{1}{2}$, i.e., $p>1-p$, so for any $v_{1}\neq v_{2}$, we have

### v3

@@ Abstract
Tensor-valued data, increasingly common in distributed big data applications like autonomous driving and smart healthcare, poses unique challenges for privacy protection due to its multidimensional structure and the risk of losing critical structural information. Traditional local differential privacy methods, designed for scalars and matrices, are insufficient for tensors, as they fail to preserve essential relationships among tensor elements. We propose TLDP, a novel LDP algorithm for Tensors that randomly decides whether to inject noise into each component with a certain probability. Such randomness substitutes for part of the perturbation noise and safeguards the raw data more effectively under the condition of guaranteeing differential privacy. To strike a better balance between utility and privacy, we incorporate a weight matrix that selectively protects sensitive regions. Both theoretical analysis and empirical findings from real-world datasets show that TLDP achieves superior utility while preserving privacy, making it a robust solution for high-dimensional tensor data.
keywords:
Local differential privacy, tensor privacy, distributed computing systems
[…]
1 Introduction
High-dimensional tensor data are passionately required in a diverse array of distributed big data applications, such as autonomous driving [1], smart healthcare [2], and face attribute recognition technology [3]. In these systems, tensor data that contains sensitive private information, such as location data, facial images, or medical records, is frequently transmitted between different parties, which creates a significant risk of privacy leakage. For example, through the intercepted tensor-valued features or model parameters (in federated systems), an attacker can easily infer whether a data type was used for training the model. Typically, Local Differential Privacy (LDP) constrains an attacker’s ability to infer details about individual data from the shared data.
@@ 1 Introduction
Developing schemes for tensors to achieve local differential privacy is a very challenging task due to the complexity of high-dimensional data. Generally, the noise introduced to the raw data is deeply affected by the sensitivity of the query function, one key factor of which is the size of the tensor. Essentially, the larger the specifications of the data tensors, the more substantial the noise tends to be, potentially leading to a decrease in the accuracy of the results. Extensive research explored LDP schemes for scalars [4, 5, 6], vectors [7, 8], and matrices [9, 10] by either adding random noise or randomly perturbing individual inputs to achieve local differential privacy and achieved promising results, but an LDP approach applicable to tensor-valued data in distributed computing systems has yet to be established. Although a few research studies like [11] introduce an LDP for tensors for distribution estimation, they cannot be directly applied for locally differential privacy protection in distributed computing systems. It is because although the perturbed tensor-valued data still retain certain statistical properties, the model analysis results on such data will be seriously affected. Therefore, in this work, we aim to tackle this challenge by developing a local differential privacy mechanism that ensures the effectiveness of the subsequent data analytics despite the high dimensionality of tensors.
@@
Both privacy and utility play a critical role when practitioners apply LDP in practical distributed computing systems. However, designing an LDP that balances the privacy and utility of data presents a significant challenge due to the inherent trade-off between the two. This issue is even more pronounced with tensors, where excessive noise or other perturbations are usually added due to the high dimensionality of tensors, which can severely diminish their usefulness. To mitigate such degradation, various methods have been explored, as seen in [12, 13, 14, 15]. However, most of them are for scalars or vectors instead of tensors, lacking a utility guarantee for tensors in distributed computing systems. In contrast, our approach ensures practical utility while satisfying $\epsilon$-LDP in distributed computing systems.
@@
In this paper, we formalize the definition of LDP for tensor-valued data and innovatively introduce a Local Differential Privacy algorithm designed for Tensors of varying dimensions named TLDP, being able to preserve the tensor-valued data’s structural information. TLDP incorporates locally differentially private and tensor-shaped noise into the tensors by randomly selecting components of the tensor with a randomized response mechanism to accomplish local differential privacy. The key idea behind it is to randomly choose whether or not to add Laplace/Gaussian noise to the tensor data so that the total amount of noise added to the tensors is lower than directly adding noise to each entity of the tensor without breaking the condition of satisfying $\epsilon$-LDP. We rigorously prove that TLDP satisfies local differential privacy both when using the Laplace/Gaussian noise and has a lower expected error than previous studies.
@@
Additionally, it is easy to observe that different parts of a tensor may have varying sensitivities in practical applications. For example, in a road scene image, the facial or license plate number information is more critical than the background for scene understanding tasks in autonomous driving applications. The data privacy of such tasks can be further improved by incorporating a precisely tailored weight matrix that indicates the areas of the tensor that are more important than others. Therefore, we design an optimized local differential privacy mechanism based on TLDP by applying varying weights to different parts of the tensor.
@@
  • 1.
    We propose an algorithm that ensures $\epsilon$-LDP for tensor-valued data named TLDP. TLDP has a lower expected error and greater utility compared to prior methods, a property independent of the tensor’s shape. To the best of our knowledge, it is the first work that addresses local differential privacy of tensor-valued data in machine learning-based distributed computing systems.
  • 2.
@@ 2 Related Work
2.1 Differential Privacy for Scalar Data
2.1.1 Differential privacy by adding random noise
In differential privacy, the Gaussian mechanism [16] and the Laplace mechanism [17] are frequently employed as privacy-preserving mechanisms. Although they do not directly apply to tensor-valued queries, our research is nevertheless grounded in the fundamental principle of preserving privacy through the injection of noise. The Gaussian mechanism introduced independent and identically distributed (i.i.d.) Gaussian noise, calibrated according to the $l_{2}$-sensitivity of the query, guarantees $(\epsilon,\delta)$-differential privacy. Similarly, by incorporating noise derived from the Laplace distribution, scaled according to the query function’s $l_{1}$-sensitivity, the Laplace mechanism enforced strong $\epsilon$-differential privacy.
Beyond these fundamental mechanisms, several sophisticated approaches have also been introduced. Mechanisms that derive their privacy guarantees from the above mechanisms are well-established, as exemplified by the composition schemes referenced in [18, 19, 20, 21]. Among these, Kairouz et al. [18] proposed an optimal composition scheme that is applicable to a broader range of noise distributions. The studies in [19, 20] provided dynamic accounting methods that adjust based on the algorithm’s runtime convergence. Moreover, Lecuyer et al. [21] ensured a global differential privacy guarantee in environments where datasets are continuously expanding.
2.1.2 Local differential privacy by randomized response to a query
@@
Warner et al. [4] introduced the randomized response mechanism in 1965, which is a technique specifically designed to preserve respondent privacy, commonly referred to as $w$-$R$. The key idea is to introduce randomness into the response process, allowing respondents to answer sensitive questions without fear of exposure, while still enabling researchers to estimate the true distribution of responses. The $w$-$R$ technique is specifically designed for discrete datasets with exactly two distinct values. To extend its applicability, two primary paths for improvement can be explored. The first involves encoding and transforming the variable’s multiple values, as demonstrated in methods such as Randomized Aggregatable Privacy-Preserving Ordinal Response (RAPPOR) [5] and S-Hist [22], to ensure compatibility with the binary nature expected by $w$-$R$. The RAPPOR employs Bloom filters and hash functions to encode a single element into a vector, then perturbs the original value through both permanent and instantaneous randomized responses. It represents a method for collecting statistical data from end-user applications while preserving anonymity and ensuring robust privacy protections. However, RAPPOR has two main drawbacks: (1) the transmission overhead between the user and the data aggregator is relatively high, as each user is required to send a vector whose length is determined by the size of its bloom filter; (2) the data collector must pre-collect a list of candidate strings for frequency counting.
@@
To address the first issue, the S-Hist method [22] offered a solution in which each user encodes the string, randomly selects a bit, applies random response techniques to perturb it, and then sends it to the data collector. This approach significantly reduces the transmission cost by minimizing the amount of data each user needs to send. For the second problem, building upon the RAPPOR-based encoding-decoding framework, Kairouz et al. [23] further proposed the $O$-$R$ method by introducing hash mapping and grouping operations. Hash mapping allows the method to focus on the encoded values rather than the original strings, eliminating the need to pre-collect a list of candidate strings. Additionally, the use of grouping operations reduces the probability of hash mapping value collisions. This combination of techniques enhances the robustness and flexibility of differential privacy in various data collection scenarios. To overcome the second limitation of $w$-$R$, one approach is to improve the distribution of the $w$-$R$ technique, making it applicable to variables with more than two values. This has been achieved through mechanisms such as $k$-$R$ [18]. In the $k$-$R$ method, the original data is preserved with probability $p$ and is flipped to any other value with a probability of $\frac{1-p}{n-1}$, where $n$ is the number of possible values. This mechanism ensures that the perturbation is uniformly distributed across all possible values, thereby preserving the privacy of the data.
2.2 Differential Privacy for Non-scalar Data
Recently, as the application of differential privacy has expanded to more scenarios, differential privacy methods designed for non-scalar datasets, i.e., vectors, matrices, and tensors, have been proposed. For instance, Li et al. [7] proposed a new privacy budget recovery mechanism, called ChainDP, which adds noise sequentially, so that the noise added by the previous user can be used by subsequent users, thereby reducing the overall noise level and estimation error. The MVG mechanism [24] is tailored for matrix-valued queries and incorporates matrix-valued noise to ensure $(\epsilon,\delta)$-differential privacy. The $l_{2}$-sensitivity of the MVG mechanism is measured using the Frobenius norm, which quantifies the discrepancy between two neighboring matrices. The IDN mechanism [25] achieved differential privacy by appending a noise tensor of equal magnitude to the data tensor. Abadi et al. [26] presented an innovative accounting approach for the Gaussian mechanism that minimizes the total additive noise while preserving the same level of privacy protection. Moreover, in order to obtain more utility under the same privacy budget, Zhang et al. [27] proposed an adaptive gradient perturbation mechanism, injecting different amount of noise according to the order of hidden layers, and Tan et al. [28] used a diagonal Hessian Matrix to generate a more effective clipping threshold of gradients.
However, LDP demands a higher level of privacy than traditional differential privacy, necessitating more noise to achieve comparable privacy guarantees, which can significantly degrade data utility due to extensive perturbation; thus additional mechanisms need to be introduced to solve the issue. Qiu et al. [8] introduced a sparse weight matrix that computes the non-zero elements of the matrix, thereby reducing the computational load of LDP when applied to large-scale datasets. But neither of them is specific to tensors. Wang et al. [10] applied differential privacy to high-order, high-dimensional sparse tensors within the IoT transmission context through individual randomized responses. But when the central server is required to receive tensor information directly from users, this approach risks disclosing the user’s original information with a certain probability, thereby undermining the purpose of privacy protection. Ren et al. [11] proposed a locally privacy-preserving scheme for crowdsensing systems to collect and build high-dimensional data from distributed users. It is improved based on the RAPPOR algorithm, which uses a hash function to convert raw data into fixed-length binary feature vectors. Due to the irreversibility of hash functions (i.e., the inability to reversely derive original attribute values from the bitstring of a Bloom filter), this method is not applicable to the scenarios in this paper where raw dataset elements are required and the query function is the identity function.
2.3 Learning With Differential Privacy
@@
In summary, we consider our paper introduces a primitive local differential privacy scheme, which fits well in distributed computing systems but certainly can be applied in any other machine learning systems, where individual users’ tensor-valued data is private and requires preservation.
3 Preliminaries
In this section, we aim to provide our readers with foundational knowledge of differential privacy to facilitate the comprehension of our work.
3.1 Local Differential Privacy
To address the scenario where the central server is untrustworthy, local differential privacy (LDP) enables data collection and analysis while ensuring strong privacy protection by perturbing data before it is shared with a central aggregator. This ensures that even the data collector cannot determine an individual’s true value with certainty. The privacy guarantee of LDP is expressed by bounding the logarithmic ratio of the output probabilities for any two possible inputs from a single user, ensuring that the private mechanism $M$ does not reveal too much about the user’s true input. This local and individualized privacy guarantee makes LDP particularly suitable for scenarios where users do not trust the data collector and need to protect their data before sharing it. Formally, given any two tensor-valued inputs from the same user $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, $y$ be the output and $M$ be the private scheme, the following condition holds:
@@ 3.1 Local Differential Privacy
$\forall y\in Range(\textbf{M}),Pr[\textbf{M}(\mathcal{X}_{1})=y]\leq e^{\epsilon}\cdot Pr[\textbf{M}(\mathcal{X}_{2})=y]$
@@
The parameter $\epsilon$ is the privacy budget of the protection mechanism M, representing the protection level of data. Generally speaking, a lower budget $\epsilon$ implies a higher level of privacy protection but usually lower data utility. So the core of research on DP and LDP lies in how to enhance data utility on the premise of ensuring privacy protection, thereby striking a balance between privacy and utility.
Problem Definition and Notations. We suppose that in this article, each user has a number of tensor-valued data, and $\mathcal{X}\in\mathbf{R}^{I_{1}\times I_{2}\times\dots\times I_{N}}$ is any one of them. We utilize $\Delta$ to represent the range of the values of a tensor in a dataset and $\Delta]$ to denote all possible values of the tensor. For example, $\Delta=256$ and $\Delta]=\{0,1,2,\dots,255\}$ for an image of the MNIST dataset, because that range of values of an image of the MNIST dataset is 0-255.Without compromising generality, we assume the input domain is $\Delta]$, i.e., $\mathcal{X}_{i_{1}i_{2}\dots i_{N}}\in[\Delta]$. If the tensor takes continuous values, it is recommended to discretize it initially. The notations used in this paper are summarized in Table 1 for easy reference.
Table 1: Notations
@@
II Product of the tensor’s dimensions, i.e., $I=I_{1}I_{2}\cdots I_{N}$
@@
In the context of LDP application, privacy protection is moved downstream from central servers to data-generating endpoints, such as users or edge devices. This shift eliminates the need for a trusted third party to handle raw data, as sensitive information is processed locally before any transmission or sharing. Therefore, our goal is to perturb the original private tensor on the client-side, ensuring that users’ sensitive information can be properly preserved without compromising data utility. To access the level of privacy protection, the parameter $\epsilon$, defined as the upper bound of the KL-divergence between the data distributions before and after perturbation, quantifies the effectiveness of privacy-preserving algorithms. As stated in Definition 1, a lower value of $\epsilon$ corresponds to stronger privacy guarantees. To access the utility, we evaluate the F1-score of the model parameters trained from the perturbed data as a measure of the data’s effectiveness. Specifically, we train the model with the perturbed data and then validate it with the unperturbed data. The closer the validation results are to the original, the higher the utility. It is evident that as privacy protection increases (i.e., $\epsilon$ decreases), the accuracy of the trained model tends to decline. Our research seeks to strike a balance between data utility and privacy protection.
3.2 Randomized Response
Randomized response is a statistical technique introduced in the 1960s by [4] for gathering data on sensitive topics while allowing respondents to maintain confidentiality. Surveys employing this method enable the accurate calculation of precise population statistics while ensuring individual privacy. Today, it is the dominant perturbation mechanism used in local differential privacy protection of scalars. For binary-valued problems, the randomized response method retains the original data with probability $p$ and flips it with probability 1p1-p. In most cases, $p>1/2$, i.e., p>1pp>1-p, so for any $v_{1}\neq v_{2}$, we have